OpenStudy (jchick):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form.

OpenStudy (jchick):

please show all work

OpenStudy (anonymous):

remember that point-slope form is y – y1 = m(x – x1). so how would you put your point and the slope in that equation?

OpenStudy (jchick):

Idk how to do these things

OpenStudy (jchick):

I know the equations and methods to find it but I don't really know how to implement it

OpenStudy (jchick):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (jchick):

That is the whole thing

OpenStudy (anonymous):

remember that m is slope, so you can substitute your slope(2) in for m. so now you have this- y - y1 = 2(x - x1). then you put in your point- so put in the 3 for 'x1' and the -2 for 'y1'. this gets you to the equation y -(-2) = 2(x - 3).

OpenStudy (anonymous):

now you just isolate y on the left side of the equal sign, and by doing so, you get slope intercept form. do you know how to isolate y?

OpenStudy (anonymous):

I really dont know how else to explain it without giving you a direct answer. @AnswerMyQuestions , could you help me explain?

OpenStudy (anonymous):

I don't understand...

OpenStudy (jchick):

I know the answer but I need to put it all into steps

OpenStudy (jchick):

I got to here but can't think what the next step is

OpenStudy (jchick):

y = 2x - 6 - 2

OpenStudy (jchick):

does it equal y = 2x - 8?

OpenStudy (jchick):

Part. 1 We know that the formula for point-slope for is y - y1 = m(x-x1) therefore (x1, y1,) = 3, 2 How is that for part one does that answer it all the way?